IST Meeting

2015-02-09 11:00
America/New York
Monday, 9 February 2015
, at 16:00 (GMT), duration : 00:00

Title: STAR Weekly IST Meeting
Community: STAR
Password: hft-ist
Meeting type: Open Meeting (Round Table)
Meeting Access Information:
SeeVoghRN Application
WebStart (Java)   
Mobile App :  Meeting ID: 101 1182   or  Link:
- Phone Bridge
 ID: 101 1182
 Password: 4774


1. Operations, Alarm Test Report - Xiaozhi
2. DB Update and Run14 Data QA - Yaping
3. Run15 Chip Status - Bingchu
4. A.O.B.
5. Minutes - Zhenyu

Attendance: Bingchu, Xiaozhi, Zhenyu (Yaping sent slides by email)

1. Operations, IST Alarm Test (Xiaozhi)
1) No issue on IST operations, first collisional data taken with IST today 9-11am today.
2) Went through the IST online QA plots together with Bingchu and Zhenyu. 
3) IST power supply and cooling alarm tests successful. A test report posted.

Action items: 
1) Make sure to understand all the QA plots. 
2) Complete the alarm test report with when the alarms are cleared during the test procedure.

2. Run15 Ist Chip Status in Online QA (BIngchu)
1) Chip 6/1/14 and 6/4/14 show as hot chips. Checked the data in pedestal and cosmic runs. Wonder why both IST/zs and IST/adc data chunks exist in st_adc_ files.

Action items:
1) ask Tonko about st_adc_ files - Answer is yes. Both chunks exist
2) plot raw adc distritributions in cosmic data and compare with those in pedestal runs

3. Run15 database and Run14 Data QA (Yaping)
1) Run15 databases are initialized, 6 pedNoise tables uploaded.
2) Run14 pre-production QA plots made. 

Action items:
1) understand why sensors 49-54 have zero hits in day 166 and 167 in Run14
2) understand fluctuations of some ladders in online QA plots in Run14